OpenStudy (anonymous):

1+4/6 divided by 1+3/6

OpenStudy (anonymous):

i think d ansr is 5

OpenStudy (angela210793):

6/6+4/6=10/6 6/6+3/6=9/6 so (10/6)/(9/6)=(10/6)*(6/9)=10/9

OpenStudy (anonymous):

first solve the numerator. This the addition of the rational number so make the denominators equal. 1+4/6 = (6+4)/6 = 10/6 On solving the denominator you get 9/6 Take it from here